OpenStudy (anonymous):

Given c is some constant Why is the derivative for d/dx Log [c x] = 1/x ? And not c Log [ c x] ?

OpenStudy (anonymous):

I mean what kind of logical thought process should I have to lock that in ?

OpenStudy (australopithecus):

Assuming it is log based 10 \[\frac{d}{dx} \log_a(u(x)) = \frac{u'(x)}{\ln(a)*u(x)}\]

OpenStudy (anonymous):

I get that d/dx Log [x] = 1/x and normally c = 0 but then if I think chain rule .. I cook up .. Log [c x] (c)(x) (c) by constant multiple c Log[ c x ] But when I put Log [c x] into mathematica, it spits out 1/x ???

OpenStudy (australopithecus):

its ln(x)?

OpenStudy (anonymous):

oh assuming oh assuming base E

OpenStudy (australopithecus):

this formula still works for ln(x) because ln(e) = 1

OpenStudy (australopithecus):

The best way two approach derivatives is to split a function into smaller functions. For example for a function f(x) = ln(ax^2) notice f(x) is just two functions g(x) = ln(x) m(x) =ax^2 differentiate m(x) an g(x), we get g(x) = ln(x) g'(x) = 1/x m(x) =ax^2 m'(x) = 2ax now plug them into the chain rule formula f'(x) = g'(m(x))*m'(x) so pluging it in f'(x) = (1/(ax^2))(2ax)

OpenStudy (australopithecus):

the constant cancels out

OpenStudy (australopithecus):

For the derivative

OpenStudy (phi):

\[ \frac{d}{du} \ln u = \frac{1}{u} du \] if u = cx then \[ \frac{d}{dx} \ln c x = \frac{1}{cx} \frac{d}{dx} cx = \frac{c}{cx} = \frac{1}{x}\]

OpenStudy (australopithecus):

remember the derivative of ax is a

OpenStudy (australopithecus):

d/dx ax = ax^0 = a(1) = a

OpenStudy (australopithecus):

and by the chain rule formula g'(m(x))*m'(x) you should see that a wont leave the function and since it is a log it will cancel out

OpenStudy (anonymous):

ahh.. gotcha. *Click*

OpenStudy (anonymous):

so my mistake was not taking the derivative for both sides, when applying the chain rule. Ln [ c x] u=cx Ln [u] (cx) 1/u c taking derivatives of both terms expand u back again c/cx 1/x simplified.

OpenStudy (australopithecus):

essentially

OpenStudy (anonymous):

probably not the best way to notate it.. but as a general identity I can say then Ln [c x] = 1/x by chain rule?

OpenStudy (anonymous):

d/dx Ln[c x] = 1/x by chain rule...

OpenStudy (australopithecus):

yes

OpenStudy (australopithecus):

Look at the two graphs, note that a constant doesnt change the slope

OpenStudy (australopithecus):

or rather a constant inside a logarithm doesn't effect the slope o

OpenStudy (australopithecus):

It just shifts the function

OpenStudy (perl):

*